# A Hierarchical Approach for Assessing the Vulnerability of Tree-Based Classification Models to Membership Inference Attack

### v1

@@ 3.4 Datasets
3.5 Experiments with Risk Correlation
It is here hypothesised that while the absolute risk cannot be predicted for a given hyperparameter combination because it depends on characteristics of the dataset, the relative risk may generalise across datasets.
Since many different risk metrics have been proposed, it is first necessary to (i) select a single representative measure to analyse the relative risk; and (ii) establish that this metric is likely to generalise across different datasets. Therefore, for each target model algorithm, pair-wise rank correlation statistics are computed to compare different MIA and target model metrics across a range of hyperparameter combinations for each dataset. Based on these results a single representative measure is selected. Then, to test whether the disclosure risk of hyperparameter combinations are likely to generalise across datasets using this metric, the results for each dataset are sorted by order of hyperparameter combination and Kendall rank correlation $\tau$ coefficients are computed to perform pair-wise comparisons.
The hypothesis is then tested by assessing the performance of models trained to predict whether a given hyperparameter combination is relatively high or low risk as a binary classification task. That is, for each target model type, a Decision Tree classifier is used to learn a simple human-interpretable ruleset that predicts whether a hyperparameter combination is relatively high or low risk. Risk is here defined as high if the previously chosen metric is in the top 20% of all hyperparameter combinations. While this is a somewhat arbitrary threshold, the aim is to prove that a simple set of rules can be identified that predict whether a hyperparameter combination poses unnecessarily high risk.
Each of these trees is thus fitted using the hyperparameter values and corresponding (discretised) MIA scores of a single dataset (mimic2-iaccd) and the extracted rules are then evaluated on the remaining datasets to test generalisation performance. In order to extract rules that are easy to interpret, the trees are restricted to a maximum depth of 5; using a simple grid search for the other hyperparameters. The best performance was found with a minimum number of samples per split of 2, a minimum number of samples per leaf of 2, the entropy criterion, and a balanced class weight.
3.6 Experiments with Structural Metrics
For each of the combinations of algorithm-dataset-hyperparameter we analysed the model produced and its predictions for the training set, and used these to capture the structural metrics described in Section 2.1.2.
We then join these with the various metrics captured from the LiRA attacks and use this combined data to examine two complementary hypotheses:
  • Hypothesis 1: Structural risks are necessary conditions for model vulnerability. In other words, if a given structural metric is 0 then there will be no MIA risk. In the following analysis we insist on 100% support for hypothesis 1 because this is a statement about safety.
  • Hypothesis 2: The presence of structural risks are sufficient conditions for model vulnerability. In other words, if (the product of a combination of) structural metric(s) is 1 there will be non-zero LiRA risk. We insist on high, but not 100% support for hypothesis 2 because this is a statement about risk.
@@ 4 Results
4.1 Results for Risk Correlation
@@ 4.1.1 Risk Correlation Across Metrics
Table LABEL:table:kendall_mimic shows the Kendall $\tau$ rank correlation statistics for each metric with pair-wise comparisons for the XGBoost classifier averaged over all datasets. Strong correlations between all of the MIA metrics can be seen. Target model generalisation error also positively correlates with risk, however target model test accuracy does not, indicating that it is strongly dataset dependant. These correlations are also seen on the other datasets with both Random Forest and Decision Tree classifiers (not shown). Since these metrics are highly correlated, the MIA AUC is used for simplicity in the remainder of the analysis. Regardless of the chosen metric, the following results should hold.
@@ 4.1.2 Risk Correlation Across Datasets
Table LABEL:table:kendall_combined shows the correlation of risk across different datasets. That is, the Kendall $\tau$ rank correlation statistics for dataset pair-wise comparisons of the MIA AUC for each target model hyperparameter combination. As can be seen, there is a strong correlation of risk with respect to hyperparameter combination across the different datasets for each of the target models. For Decision Tree models the mean is 0.64 (min=0.45, max=0.81); for Random Forest models the mean is 0.72 (min=0.51, max=0.88); and for XGBoost models the mean is 0.67 (min=0.48, max=0.84). This suggests that it may be possible to train a classification model that identifies the relative risk of a hyperparameter combination for most datasets.
@@ 4.1.3 Risk Prediction Across Datasets
Table LABEL:table:accuracy_combined shows the precision, recall, and accuracy of the simple rulesets at predicting whether a given target model hyperparameter combination is in the top 20% highest MIA AUC (high risk) or in the bottom 80% (low risk). Note that the risk prediction rules for each target model were derived by training a Decision Tree classifier on the results from the mimic2-iaccd dataset then distilling the tree into a set of rules. The mean accuracy on the unseen datasets is 91.4% (min=89.1%, max=93.9%) for Decision Tree target models; 95.2% (min=92.9%, max=96.2%) for Random Forest; and 94.1% (min=90.8%, max=97.2%) for XGBoost target models, showing that the very simple ruleset is able to accurately differentiate the hyperparameter combinations with the highest relative risk on each dataset.
@@
Appendix A.1 shows the extracted and compressed rules for identifying high risk hyperparameter combinations in tree-based models. As expected, model complexity is a significant risk factor with larger values for max_depth and n_estimators being key features along with smaller values of min_samples_leaf and min_child_weight. A deeper tree with fewer samples per node allows the model to learn more intricate relationships in the data, potentially overfitting. Similarly, a larger number of trees in an ensemble can lead to overfitting if not carefully managed.
@@
A combination of higher potential model complexity and greedy search algorithms are therefore much more likely to lead to models that are highly susceptible to overfitting and thus pose a higher disclosure risk in practice.
@@ 4.2 Results for Structural Risks
To examine support for the two different hypotheses stated in Section 3.6 we aggregated the results from LiRA and structural attacks on all (73610) combinations of dataset-algorithm-hyperparameters.
@@
Hypothesis 1 proposed that when there was no structural risk, the LiRA attack risk metrics would also be near zero. The second observation above is illustrated in Figure LABEL:fig:heatmaps (left). This shows conclusive evidence that this hypothesis can be rejected. In other words, structural risk is not necessary for vulnerability to membership inference.
Hypothesis 2 took the alternative position that the presence of structural risk was predictive of a model being vulnerable to a LiRA attack. As Figure LABEL:fig:heatmaps (right) shows there are only a handful of such cases that do not also display MIA vulnerability.
While these counterexamples cannot be ignored, from the perspective of an analyst making pragmatic decisions about whether a trained model poses a privacy risk, there is a clear message. The combined structural measure presents sufficient but not necessary evidence of MIA vulnerability, without the need to run computationally expensive attacks.
4.2.1 Analysis of Where LiRA Attack Metrics are Statistically Significant
@@ 5 Conclusion
This article has presented the first empirical analysis of tree-based classification model hyperparameters in terms of their susceptibility to MIA. It has been shown that simple human-interpretable rulesets can predict with approximately 90–95% accuracy whether a given hyperparameter combination is one of the 20% most vulnerable. Significantly, it has been shown that for most datasets, high target model accuracy is still obtainable with hyperparameter combinations that are not in this highest risk category.
The use of more complex rules/models would likely increase the accuracy of risk prediction, however the use of a simple Decision Tree to extract the ruleset enables easy interpretation by researchers so that they can more easily adjust their preferred hyperparameters into a safer region that maintains sufficient classification accuracy.
It is important to note that these rules cannot determine whether a hyperparameter combination can be considered safe since the absolute risk scores are highly dependant on the dataset. However, these results suggest that the relative risk of hyperparameters is consistent across most datasets and that these highest risk values can reasonably be removed from the hyperparameter tuning process and consequently before releasing a trained model into the wild without a significant loss of model performance. Some datasets/applications may still require some privacy trade-off for sufficient accuracy, but it seems reasonable that as the first stage of risk assessment it should be incumbent on the researchers wishing to release a model to first show that it is necessary to use these highest risk hyperparameters.
@@
Current work is integrating the learned rulesets and structural attacks within the open source sacro-ml [35] Python package to provide a comprehensive hierarchical framework to automatically perform these risk assessments in addition to running current state-of-the-art MIAs. Future work may explore additional target models, hyperparameters, and datasets. Including, for example, the consideration of neural networks.
@@ Acknowledgements
This work was funded by UK Research and Innovation (Grant Numbers MC_PC_23006 and MC_PC_21033) as part of Phase 1 of the Data and Analytics Research Environments UK (DARE UK) programme, delivered in partnership with Health Data Research UK (HDR UK) and Administrative Data Research UK (ADR UK).

### v2

@@ 4 Results
4.1 Results for Model Hyperparameter Risks
@@ 4.1.1 Risk Correlation Across Metrics
Table LABEL:table:kendall_mimic shows the Kendall $\tau$ rank correlation statistics for each metric with pair-wise comparisons for the XGBoost classifier averaged over all datasets. Strong correlations between all of the MIA metrics can be seen. Target model generalisation error also positively correlates with risk, however target model test accuracy does not, indicating that it is strongly dataset dependant. These correlations are also seen on the other datasets with both Random Forest and Decision Tree classifiers (not shown). Since these metrics are highly correlated, this provides strong evidence in support of hypothesis 1.1 and the MIA AUC is used for simplicity in the remainder of the analysis. Regardless of the chosen metric, the following results should hold.
@@ 4.1.2 Risk Correlation Across Datasets
Table LABEL:table:kendall_combined shows the correlation of risk across different datasets to test hypothesis 1.2. That is, the Kendall $\tau$ rank correlation statistics for dataset pair-wise comparisons of the MIA AUC for each target model hyperparameter combination. As can be seen, there is a strong correlation of risk with respect to hyperparameter combination across the different datasets for each of the target model types. For Decision Tree models the mean is 0.64 (min=0.45, max=0.81); for Random Forest models the mean is 0.72 (min=0.51, max=0.88); and for XGBoost models the mean is 0.67 (min=0.48, max=0.84). This suggests that it may be possible to train a classification model that identifies the relative risk of a hyperparameter combination for most datasets.
@@ 4.1.3 Risk Prediction Across Datasets
Table LABEL:table:accuracy_combined shows the precision, recall, and accuracy of the simple rulesets at predicting whether a given target model hyperparameter combination is in the top 20% highest MIA AUC (high risk) or in the bottom 80% (low risk). Note that the risk prediction rules for each target model were derived by training a Decision Tree classifier on the results from the mimic2-iaccd dataset then distilling the tree into a set of rules. The mean accuracy on the unseen datasets is 91.4% (min=89.1%, max=93.9%) for Decision Tree target models; 95.2% (min=92.9%, max=96.2%) for Random Forest; and 94.1% (min=90.8%, max=97.2%) for XGBoost target models, showing that the very simple ruleset is able to accurately differentiate the hyperparameter combinations with the highest relative risk on each dataset, supporting hypothesis 1.3.
@@
Beyond overfitting, MIA vulnerability is significantly influenced by model type and structure [45]. This is confirmed here with the rules learned from the mimic dataset for identifying high risk hyperparameter combinations in tree-based models. As might be expected, these rules consistently associate ‘high risk’ MIA vulnerability with model configurations that promote higher complexity. For example, across all model types (Decision Tree, Random Forest, and XGBoost), a greater max_depth is a recurring condition for high risk. Similarly, a larger number of n_estimators in ensemble methods (Random Forest and XGBoost) also points to increased vulnerability.
However, the rules reveal nuances beyond sheer size. Specific hyperparameter choices that dictate the model’s learning capacity and flexibility are critical:
  • For Decision Trees, risk factors include: utilising splitter=best rather than random; allowing small values for min_samples_leaf or min_samples_split; or considering all features for splits (max_features=None).
  • In Random Forests, conditions such as bootstrap=False (where each tree is trained on the entire dataset, potentially reducing diversity) or specific min_samples_split thresholds when combined with other hyperparameters contribute to high risk.
  • For XGBoost, a low min_child_weight, which permits the creation of tree nodes with fewer samples, is a key indicator of increased vulnerability.
These specific hyperparameter settings effectively enable the models to create more intricate and numerous decision boundaries. This increased capacity to fit the training data closely not only reflects a higher Vapnik-Chervonenkis (VC) dimension (hence the link to generalisation error) but also, as the rules suggest, heightens the model’s susceptibility to MIA.
@@
A combination of higher potential model complexity and greedy search algorithms are therefore much more likely to lead to models that are highly susceptible to overfitting and thus pose a higher disclosure risk in practice. The risk prediction rules may be found in Appendix A.1.
@@ 4.2 Results for Structural Risks
To examine support for the two different hypotheses stated in Section 3.2 we aggregated the results from LiRA and structural attacks on all (73610) combinations of dataset-algorithm-hyperparameters. We also created a Combined_AND structural risk as the logical conjunction (AND) of the degrees of freedom (DoF), $k$-anonymity (K_anon) and class disclosure (lowvals_cd) structural measures.
@@
Hypothesis 2.1 proposed that when there was no structural risk, the LiRA risk metrics would also be near zero. The second observation above is illustrated in Figure LABEL:fig:heatmaps (left). This shows conclusive evidence that this hypothesis can be rejected. In other words, structural risk is not necessary for vulnerability to MIA, at least in the form in which structural risk is currently quantified.
Hypothesis 2.2 took the alternative position that the presence of structural risk was predictive of a model being vulnerable to LiRA. As Figure LABEL:fig:heatmaps (right) shows there are only a handful of such cases that do not also display MIA vulnerability.
While these counterexamples cannot be ignored; from the perspective of an analyst making pragmatic decisions about whether a trained model poses a privacy risk, there is a clear message. The combined structural measure, Combined_AND presents sufficient but not necessary evidence of MIA vulnerability, without the need to run computationally expensive attacks. A risk prediction classifier based on this simple rule has:
  • high precision (0.927–0.993, depending on the specific MIA risk metric being predicted)
  • but relatively low recall (0.461–0.514, depending on the attack metric).
It is worth reiterating that this is only a filtering approach. A prediction of ‘not unsafe’ should not be interpreted as meaning the target is safe, rather that further investigation is needed (because the AND rule has low recall). This is somewhat akin to the distinction between ‘innocent’, ‘guilty’, and ‘not proven’ in some legal jurisdictions.
4.2.1 Analysis of Where MIA Metrics are Statistically Significant
@@ 5 Conclusion
This article has presented the first empirical analysis of tree-based classification model hyperparameters in terms of their susceptibility to MIA. Surprisingly, regardless of MIA success metric, the rank order correlation of disclosure risk is seen to be preserved across datasets. It has been shown that simple human-interpretable rulesets learned on one dataset can predict with approximately 90–95% accuracy whether a given hyperparameter combination is one of the 20% most vulnerable on other unseen datasets. Significantly, it has been shown that for most datasets, high target model accuracy is still obtainable with hyperparameter combinations that are not in this highest risk category.
The use of more complex rules/models such as deep neural networks would likely increase the accuracy of risk prediction, however the use of a simple Decision Tree to extract the ruleset enables easy interpretation by researchers so that they can more easily adjust their preferred hyperparameters into a safer region that maintains sufficient classification accuracy. Future work may explore the use of rule-based learners such as the XCSF learning classifier system [34] to extract interpretable rules with improved accuracy.
While these simple rules offer valuable insights, it is important to acknowledge that they cannot definitively determine ‘safe’ hyperparameter combinations, as absolute risk scores are highly dependent on dataset characteristics such as distribution, imbalances, outliers, and size. Nevertheless, our results suggest that the relative risk of hyperparameters can remain surprisingly consistent across diverse datasets. This consistency implies that hyperparameter values associated with the highest risk can reasonably be excluded from the tuning process—and consequently, before running expensive MIAs—without significant loss of performance. Although further dataset testing with the rules extracted here may reveal limitations in generalisation, the existence of dataset classes where rules learned from one dataset successfully generalise to others presents an exciting opportunity to exploit similarities within secure data facilities such as TREs, which often house large numbers of related datasets.
Some datasets/applications may still require some privacy trade-off for sufficient utility, but it seems reasonable that as the first stage of risk assessment it should be incumbent on the researchers wishing to release a model to first show that it is necessary to use these highest risk hyperparameters. In the event that the rules do not appear to be generalising to a particular dataset after sampling some hyperparameter values, then the use of surrogate modelling to efficiently identify the best privacy-utility trade-off could be used (i.e., generalising within the dataset rather than across datasets), however this may not be possible for very large models and datasets.
@@
Current work is integrating the learned rulesets and structural attacks within the open source SACRO Python toolkit [47, 33] to provide a comprehensive hierarchical framework to perform these risk assessments, in addition to running current state-of-the-art MIAs. Future work may explore additional target models, hyperparameters, and datasets. Including, for example, the consideration of non-tree models such as neural networks. Potentially, a similar approach could be undertaken by using feature encoding schemes similar to those used for the hyperparameters when surrogate modelling neural architectures, e.g., hand-crafted features such as the number of parameters, depth, width, number of convolutions, or sequence based representations such as tokenising the architecture [52, 29]. The structural metrics presented here also have similarities with zero-cost proxies for estimating the accuracy of deep neural networks [26] and issues such as overparameterising and lack of regularisation hyperparameters may be indicative of MIA risk.
The data produced from experimentation and analysis code used to produce the figures and statistics are available in [35].
@@ Acknowledgements
This work was funded by UK Research and Innovation (Grant Numbers MC_PC_23006, MC_PC_21033, and MC_PC_24038) as part of the Data and Analytics Research Environments UK (DARE UK) programme, delivered in partnership with Health Data Research UK (HDR UK) and Administrative Data Research UK (ADR UK).